### v1

# LiT: Delving into a Simplified Linear Diffusion Transformer for Image Generation

Abstract
In commonly used sub-quadratic complexity modules, linear attention benefits from simplicity and high parallelism, making it promising for image synthesis tasks. However, the architectural design and learning strategy for linear attention remain underexplored in this field. In this paper, we offer a suite of ready-to-use solutions for efficient linear diffusion Transformers. Our core contributions include: (1) Simplified Linear Attention using few heads, observing the free-lunch effect of performance without latency increase. (2) Weight inheritance from a fully pre-trained diffusion Transformer: initializing linear Transformer using pre-trained diffusion Transformer and loading all parameters except for those related to linear attention. (3) Hybrid knowledge distillation objective: using a pre-trained diffusion Transformer to help the training of the student linear Transformer, supervising not only the predicted noise but also the variance of the reverse diffusion process. These guidelines lead to our proposed Linear Diffusion Transformer (LiT), an efficient text-to-image Transformer that can be deployed offline on a laptop. Experiments show that in class-conditional 256256 and 512512 ImageNet benchmark LiT achieves highly competitive FID while reducing training steps by and compared to DiT. LiT also rivals methods based on Mamba or Gated Linear Attention. Besides, for text-to-image generation, LiT allows for the rapid synthesis of up to 1K resolution photorealistic images. Project page: https://techmonsterwang.github.io/LiT/.
1 Introduction
Diffusion models [29, 67] based on Transformer [71], such as PixArt- [7] and Stable Diffusion 3 [16], recently demonstrate potential commercial value in the field of high-resolution text-to-image generation. However, one of the core computational modules i.e., self-attention, exhibits a quadratic complexity with respect to sequence length, leading to increased latency and GPU memory usage in high-resolution scenarios, which makes it less feasible for resource-limited edge devices, as illustrated in Fig. 4 and Fig. 8 (Sec. B of the appendix).
Accordingly, existing works [69, 87] adopt sub-quadratic modules, such as State Space Models (SSM) [20, 21] and Gated Linear Attention (GLA) [78], to construct diffusion models for image generation. Their designs are illustrated in Fig. 2-(a). While they all achieve sub-quadratic complexity, the overall complexity differs substantially. First, both SSM blocks and GLA rely on recurrent states, leading to iterative computation that limits parallelization opportunities. Second, these modules have complex computational graphs and involve low-arithmetic-intensity operations (e.g., element-wise multiplication).
In contrast, linear attention [38, 3, 23, 1, 6] linear attention provides an efficient alternative to self-attention, with a simple design and easy parallelization. It can be observed that linear attention speeds up generation compared to self-attention at the same resolution. For example, at a 2048 resolution, linear attention is nearly 9 faster (shown in Fig. 4-(a) and (b)). Besides, linear attention allows for fewer GPU memory for higher-resolution image generation. For instance, the GPU memory for generating 2048px image with DiT-S/2 can be dropped from 14GB to 4GB if replacing with the linear attention counterpart. Therefore, training a high-performance diffusion Transformer with linear attention is valuable for edge-side devices with strict latency and computational resource requirements.
Nevertheless, how to effectively and rapidly train linear attention-based diffusion Transformer for challenging image generation tasks? On one hand, although linear attention successfully replicates the effectiveness of self-attention in visual recognition, architectural designs for linear attention in image generation remain an under-explored issue. On the other hand, training diffusion models from scratch is costly. For example, as statistically reported in [8], training RAPHAEL [76] requires 60K A100 GPU days. As a result, cost-effective training strategies for linear Diffusion Transformers are still worth exploring.
In this work, we systematically study the diffusion Transformer implementation by pure linear attention, from both architectural design and training strategy. For architectural refinement, we draw on successful practices from multimodal understanding models and observe the free lunch effect of the linear attention. For training strategies, we focus on two cost-effective methods—weight inheritance and knowledge distillation—to improve the training process.
Our findings about architectural refinement and optimization strategies on efficient linear DiT can be summarized as the following practical guidelines: 1) employ Simplified Linear Attention in image understanding tasks; 2) use few (e.g., 2) heads in the linear attention for better performance but no latency increase; 3) initialize the weights of the linear diffusion Transformer from a well-trained self-attention based diffusion model; 4) load all weights except for the linear attention to achieve cost-effective training; 5) adopt hybrid distillation objective—distilling not only the predicted noise but also variances of the reverse diffusion process—to facilitate the student diffusion Transformer.
| (a) Comparison of subquadratic computational complexity designs | (b) Roadmap for improving LiT |
These guidelines lead us to develop LiT, a diffusion Transformer using pure linear attention. LiT is trained in a highly cost-effective manner while maintaining a high-resolution-friendly property during inference, and can be deployed offline on a Windows 11 laptop. In class-conditional image generation on the ImageNet [13] 256256 benchmark, LiT-S/B/L (patch size of 2) with only 100K training steps outperforms corresponding 400K steps pre-trained DiT-S/B/L in terms of Fréchet Inception Distance (FID) [27]. For large-scale model on the ImageNet 256256 and 512512 benchmark, LiT-XL/2 competes comparably to DiT-XL/2 (2.32 vs. 2.27, 3.69 vs. 3.04 in FID) with only and of training steps (respectively for 256256 and 512512 settings). In challenging text-to-image generation, our LiT-0.6B can stably generate highly photorealistic images (Fig. 7) with only 40K training steps. As a result, we summarize the main contributions:
We introduce a linear diffusion Transformer, LiT, which optimizes the latency-capacity trade-off by using a few heads (e.g., 2) in linear attention.
We initialize LiT with weight inheritance to ensure a cost-effective training process, while achieving linear computational complexity that facilitates high-resolution generation scenarios during inference.
We propose a hybrid distillation approach, thereby successfully apply LiT to class-conditional image generation and text-to-image generation.
The pretrained LiT-0.6B model can be deployed on a Windows 11 laptop to generate high-resolution, photo-realistic images offline based on user prompts.
2 Related Work
Diffusion models [29, 67, 68] have proven effective in image generation, and can be implemented using Transformer architecture based on self-attention [71], e.g., DiT [53], PixArt- [8], PixArt- [7], and Lumina-T2X [18].
From an architectural perspective, recent work begins focusing on efficient diffusion models [69, 77, 34], based on Mamba [21, 11] and RWKV [54] to compete with DiT. Meanwhile, linear attention [38], as a simple linear complexity token mixer [83], is being applied in diffusion Transformers for image synthesis, e.g., Mediator [58], DiG [87] for class-conditional image generation, and Sana [75], LinFusion [46] for text-to-image generation. In summary, existing work trains diffusion Transformers with linear attention from scratch. In contrast, our LiT aims for cost-effective training by exploring linear attention design, weight inheritance and knowledge distillation. Moreover, our approach has been validated on both class-conditional image generation and text-to-image generation.
3 Preliminaries and Motivation
3.1 Preliminaries
Diffusion models.
Suppose we have a clean image , and progressively add noise at each step to obtain . The forward (diffusion) process is defined as: , where and . Diffusion models [29, 67, 48, 74] learn a network to reverse the diffusion process: , where and denote the mean and variance of the reverse process. In DDPM [29], a simple training objective is demonstrated effective in training diffusion models: , where represents noise of the step, expected to be approximated by the output of a denoising network . can be realized using Transformers [53, 49, 8, 7] with time-intensive self-attention, calling for cheap alternatives.
Linear attention.
For a standard softmax attention with heads, sequence length , and hidden dimension , denoting as query, key and value, its output can be expressed as:
where is a similarity measurement function. In linear attention, the aforementioned function is modified into a simplified form with a nonlinear kernel , i.e., . Thus, Eq. 1 can be rewritten as Eq. 2 and simplified to Eq. 3 using the associative property of multiplication.
The computational complexities of softmax attention and linear attention are and , respectively. Detailed presentation is provided in Sec. 4.2.
| (a) Latency vs resolution | (b) GPU memory vs resolution | (c) Cosine similarity among heads |
3.2 Motivation
For scenarios requiring high-resolution generation (e.g., 10241024), self-attention is latency-intensive. In a common setup, using a latent diffusion Transformer [62] with an 8 downsampling VAE [39] encoder (as in DiT-XL/2 [53]), the input latent resolution is 128128, bringing a sequence length of 4096. Accordingly, we have . For higher resolutions, this ratio will increase.
Based on this analysis, self-attention is not well-suited for commercially valuable high-resolution generation tasks. Can we use cheap linear attention to replace self-attention in diffusion models? To this end, we aim to explore how to cost-effectively train a linear DiT with photorealistic generation quality by focusing on linear attention design, weight inheritance, and diffusion knowledge distillation paradigm. We discuss our experiments and conclusions in detail below.
4 Cost-Effectively Training of Linear DiT
In this section, we provide the LiT design trajectory. Given the limited research on linear DiT, we first build a suitable linear attention baseline for generative tasks, before investigating advanced training strategies. Our baseline shares the same macro architecture as DiT [53], which has recently proven effective for text-to-image generation [8, 7, 44], with the only difference being that we replace self-attention with linear attention that is computationally efficient in high-resolution image synthesis. Additionally, we employ latent diffusion [62] consistent with DiT. LiT generally follows DiT’s configuration (for S/B/L/XL), except the number of attention heads. All experiments are conducted on the class-conditional ImageNet [13] 256256 benchmark, with models trained for 400K iterations using a batch size of 256. We report the FID-50K [27], Inception Score [65] and Precision/Recall [40] metrics. The radar chart of the experimental results is shown in Fig. 2-(b). Detailed results can be found in Sec. D of the appendix.
4.1 Simplified Linear Attention in DiT
| DiT | Attention | FID-50K () | IS () |
| S/2 | softmax | 68.40 | - |
| S/2 | ReLU linear | 88.46 | 15.11 |
| S/2 | Simplified linear (ReLU) | 63.66 | 22.16 |
| S/2 | focused linear (ReLU) | 63.05 | 22.49 |
| S/2 | focused linear (GELU) | 70.83 | 19.41 |
| B/2 | softmax | 43.47 | - |
| B/2 | ReLU linear | 56.92 | 25.80 |
| B/2 | Simplified linear (ReLU) | 42.11 | 34.60 |
| B/2 | focused linear (ReLU) | 40.58 | 35.98 |
| B/2 | focused linear (GELU) | 58.86 | 24.23 |
Guideline 1: Simplified linear attention in image recognition is sufficient for DiT-based image generation.
Linear attention has been successfully applied in perceptual ViTs [3, 23, 24, 1] as a replacement for self-attention, which are used to extract high-level features for image classification. On the other hand, pioneering works apply them in generative models [58, 75, 46], where they are used to predict low-level noise for text-to-image synthesis. Inspired by the successful practices in ViTs [15], we now investigate how to simplify the linear attention design in LiT while effectively preserving the fidelity of generated images. We start with a common practice, i.e., ReLU linear attention [38], which uses the ReLU activation as a kernel function to approximate the non-linearity of the softmax function. For performance reference, we compare it to DiT, where any performance difference can be attributed to the impact of linear attention on generation quality. As shown in Tab. 1, compared to DiT, LiT-S/2 and B/2 with a trivial ReLU linear attention leads to an unacceptable performance drop, with FID decreasing by approximately 20/13 for the S/2 and B/2 models. The results suggest that linear attention commonly used in visual recognition have room for improvement in noise prediction tasks.
We then explore the following approaches: (1) simplified linear attention [22] (ReLU linear attention with a depthwise convolution (DWC) [33, 9]; (2) focused linear attention [23]; (3) replacing the ReLU kernel with GELU [26] function, commonly used in Transformers [71]. Each of these choices maintains linear complexity, preserving LiT’s advantage in computational efficiency. We use a relatively large convolution kernel (i.e., 5) to ensure a sufficiently large receptive field when predicting noise. Tab. 1 shows that adding DWC alone improves generation quality beyond softmax attention, with FID of 63.66 and 42.11 for S/2 and B/2. We attribute the result to the model’s tendency to focus on information from neighboring pixels when predicting noise for a given pixel, thus requiring the assistance of convolution. Meanwhile, we identify that the focused function has limited effectiveness, which we attribute to its design motivation to help linear attention focus on specific regions. This feature may suit classification models, but may not be necessary for noise prediction.
Remark 1. Given that a major motivation for using linear attention is to address latency issues with high-resolution images, we use simplified linear attention for simplicity and performance. Next, we focus on heads of the linear attention for a better latency-performance trade-off.
| (a) Results of small (S) and base (B) model | (b) Results of large (L) and Xlarge (XL) model |
4.2 Linear Attention with Few Heads
Guideline 2: Using few heads in the linear attention increases computation but not latency.
In practice, DiT-S/B/L/XL are designed to have 6/12/16/16 heads, respectively. Denoting the number of heads, sequence length, and hidden dimension as , GMACs of multi-head self-attention and linear attention are and , respectively, as follows:
Theoretically, is independent of the number of heads, but in linear attention, has a negative correlation with it. Intuitively, using many heads can loose the computational stress. But on the contrary, we prefer to use few heads, based on the identified free lunch effect of linear attention. We illustrate the effect in Fig. 5, where we visualize the latency and theoretical GMACs of linear attention in DiT-S/B/L/XL models with different head numbers. We use one NVIDIA A100 GPU to generate 256256 resolution images with a batch size of 8, averaging the results over 30 experiments. The results show that decreasing the number of heads leads to a steady increase in theoretical GMACs but not the practical latency, which even shows a decrease. Accordingly, we summarize it as the free lunch effect of linear attention. We also observe this phenomenon on the standard NVIDIA V100 GPU, as reported in Sec. C of the appendix. As such, we argue that setting a small number of heads enables a high theoretical compute budget, which, according to scaling law [45, 37, 31], allows the model to reach a higher upper bound in generation performance. The argument can be empirically demonstrated in Tab. 2. For different model scales, equipping linear attention with a few heads (e.g., 2, 3, 4) outperforms the default settings in DiT. Instead, using an excessive number of heads (e.g., 96 for S/2 or 192 for B/2) seriously hinders the generation quality.
Besides, we also evaluate the necessity of using many heads in linear attention from a representation similarity lens. Specifically, we visualize the average cosine similarity among the attention maps of different heads in LiT-S/2 (with 6 heads). As depicted in Fig. 4-(c), for different denoising steps or layers, the attention across different heads is quite homogeneous, with the average cosine similarity reaching up to above 0.5. The observation suggests that a small number of heads dominate the major information in the linear attention, which positively supports the empirical generative results in Tab. 2.
| DiT | Head | FID-50K () | IS () | Prec. () | Rec. () |
| S/2 | 1 | 64.42 | 21.54 | 0.380 | 0.574 |
| S/2 | 2 | 63.24 | 22.07 | 0.385 | 0.570 |
| S/2 | 3 | 63.21 | 22.08 | 0.386 | 0.583 |
| S/2 | 6 | 63.66 | 22.16 | 0.383 | 0.580 |
| S/2 | 48 | 78.76 | 17.46 | 0.322 | 0.482 |
| S/2 | 96 | 116.00 | 11.49 | 0.224 | 0.261 |
| B/2 | 1 | 41.77 | 34.78 | 0.487 | 0.631 |
| B/2 | 2 | 41.39 | 35.59 | 0.494 | 0.631 |
| B/2 | 3 | 40.86 | 35.79 | 0.497 | 0.629 |
| B/2 | 12 | 42.11 | 34.60 | 0.484 | 0.631 |
| B/2 | 96 | 68.30 | 20.45 | 0.375 | 0.531 |
| B/2 | 192 | 112.39 | 12.07 | 0.240 | 0.282 |
| L/2 | 1 | 24.46 | 57.36 | 0.600 | 0.637 |
| L/2 | 2 | 24.37 | 57.02 | 0.599 | 0.622 |
| L/2 | 4 | 24.04 | 59.02 | 0.597 | 0.636 |
| L/2 | 16 | 25.25 | 54.67 | 0.587 | 0.632 |
| XL/2 | 1 | 21.13 | 65.06 | 0.619 | 0.632 |
| XL/2 | 2 | 20.66 | 65.39 | 0.624 | 0.636 |
| XL/2 | 4 | 20.82 | 65.52 | 0.619 | 0.632 |
| XL/2 | 16 | 21.69 | 63.06 | 0.617 | 0.628 |
4.3 Inheriting Weights from Diffusion Transformer
LiT shares some structural components with DiT, allowing weights to be inherited from the pre-trained DiT. These weights contain rich knowledge related to noise prediction, which is expected to transfer to LiT in a cost-effective way. Therefore, we load the pre-trained DiT weights into LiT, except for the linear attention. This includes parameters related to FFN, adaptive layer norm (adaLN), positional encoding, and conditional embedding. To thoroughly investigate this, we conduct ablation experiments to explore how the degree of training on inherited weights affects the model’s generation performance.
Guideline 3: Linear diffusion Transformer should be initialized from a converged DiT.
We first pre-train DiT-S/2 [53] for five types of iterations—200K, 300K, 400K, 600K, and 800K—and in separate experiments, load these pre-trained weights into LiT-S/2, keeping only the linear attention parameters randomly initialized. The initialized LiT-S/2 is then trained for 400K iterations on class-conditional ImageNet, with results shown in Tab. 3. Some interesting findings are identified: (1) DiT’s pre-trained weights, even trained only 200K steps, play a significant role, improving the FID from 63.24 to 57.84. (2) Using exponential moving average (EMA) [57] of pre-trained weights has minimal impact. (3) Longer pre-training of DiT makes them better suited as initialization for LiT, even though the architectures are not fully aligned.
Remark 3. We suppose a possible explanation for this phenomenon is the functional decoupling of different modules in diffusion Transformers, as explored in large language models for content generation [79, 73]. Basically, although DiT and LiT have different architectures, their shared components (e.g., FFN and adaLN) act quite similarly. As a result, the knowledge in the pre-trained parameters of these components can be transferred, and extensive pre-training may not hinder the optimization of the non-transfered parts.
| Load | Iter. | FFN | Modu. | Attention | FID-50K () |
| model | 400K | ✓ | ✓ | ✗ | 56.07 |
| ema | 400K | ✓ | ✓ | ✗ | 56.07 |
| model | 200K | ✓ | ✓ | ✗ | 57.84 |
| model | 300K | ✓ | ✓ | ✗ | 56.95 |
| model | 400K | ✓ | ✓ | ✗ | 56.07 |
| model | 600K | ✓ | ✓ | ✗ | 54.80 |
| model | 800K | ✓ | ✓ | ✗ | 53.83 |
| model | 600K | ✓ | ✓ | Q, K, V | 55.29 |
| model | 600K | ✓ | ✓ | K, V | 55.07 |
| model | 600K | ✓ | ✓ | V | 54.93 |
| model | 600K | ✓ | ✓ | Q | 54.82 |
| model | 600K | ✓ | ✓ | O | 54.84 |
Guideline 4: Projection matrices of query, key, value, and output in linear attention should be initialized randomly.
In LiT, some of the weights in linear attention overlap with those in self-attention of DiT, including the query, key, value, and output projection matrices. Despite the differences in computation paradigms, these weights can be directly loaded from DiT into the LiT without additional processing. However, whether this can accelerate its convergence remains an open question. We use a DiT-S/2 pre-trained for 600K iterations to conduct ablation experiments. Five different types of loading strategies are explored, including (1) query, key, and value projection matrices; (2) key and value projection matrices; (3) the value projection matrix; (4) the query projection matrix; and (5) the output projection matrix. Evaluation results are presented in Tab. 3. In comparison to the baseline without loading self-attention weights, none of the explored strategies shows better generative performance. The phenomenon can be attributed to differences in computation paradigms. Specifically, linear attention directly computes the product of key and value matrices, whereas self-attention does not. Consequently, weights related to key and value from self-attention offer limited benefit to linear attention. Clearly, as presented in Tab. 3, inheriting the weights of the key and value projector leads to a larger degradation (FID of 55.29 and 55.07).
| Iter. | Teacher | FID-50K () | IS () | ||
| 800K | DiT-S/2 | 0.1 | 0.0 | 55.11 | 26.28 |
| 800K | DiT-XL/2 | 0.0 | 0.0 | 53.83 | 27.16 |
| 800K | DiT-XL/2 | 0.1 | 0.0 | 53.05 | 27.43 |
| 800K | DiT-XL/2 | 0.05 | 0.0 | 53.41 | 27.26 |
| 800K | DiT-XL/2 | 0.5 | 0.0 | 51.13 | 28.89 |
| 800K | DiT-XL/2 | 0.1 | 0.05 | 52.76 | 27.70 |
| 800K | DiT-XL/2 | 0.0 | 0.05 | 53.49 | 27.26 |
| 800K | DiT-XL/2 | 0.05 | 0.05 | 53.14 | 27.46 |
| 800K | DiT-XL/2 | 0.5 | 0.05 | 50.79 | 29.17 |
Remark 4. Inheriting weights from a LiT-like pre-trained model is a favorable approach. We recommend loading all pre-trained parameters except for the linear attention, as it is easy to implement and well-suited for diffusion models based on the Transformer macro architecture. With substantial GPU resources, extensive pre-training can further improve linear LiT’s convergence and optimization, resulting in an efficient but strong model.
4.4 Distilling Both Predicted Noise and Variance
Guideline 5: Hybrid distillation is neccessary for student linear diffusion Transformer. We distill not only the predicted noise but also variances of the reverse diffusion process, but in a moderate way.
Knowledge distillation [28] commonly employs a teacher network to help the training of a lightweight student network, which has been validated in ConvNets [19] and ViTs [72, 70, 85]. For diffusion models, distillation typically focuses on reducing the sampling steps of the target model [81, 80, 64]. In contrast, we focus on how a heavy teacher model can aid the learning of an efficient student from an architecture perspective, while maintaining the sampling steps [46].
A pre-trained DiT-S/2/XL/2 serves as the teacher network to distill the student LiT-S/2. For strong performance, we use a DiT-S/2 pre-trained for 800K iterations for initialization, following Guideline 3&4. Denote as the noised input latent in timestep , where , , are height, width and the number of channels of the input latent. denote network parameters of the teacher diffusion Transformer and the student LiT network , respectively. As denoising diffusion models, the output of DiT and LiT include both noise predictions and variances of the reverse diffusion process . DDPM [29] fixes the variance and optimizes model parameters using . Naturally, a straightforward idea is to align the student network’s noise prediction capabilities with those of the teacher through a mean squared error (MSE) objective. Thus, we have:
where is the normalization factor. The impact of distilling the predicted noise is reported in Tab. 4. It shows that using a stronger teacher model (e.g., DiT-XL/2) improves the student model’s generation capbility (51.13 vs. 53.83, measured by FID). Moreover, increasing the weight of the distillation objective from 0.05 to 0.5 consistently boosts performance, suggesting that the cheap linear attention can benefit from the supervision of the self-attention in diffusion Transformers.
| Class-Conditional ImageNet 256256 | |||||
| Model | Training Steps | FID | IS | Prec. | Rec. |
| DiT-S/2 [53] | 400K | 68.40 | - | - | - |
| DiT-B/2 | 400K | 43.47 | - | - | - |
| DiT-L/2 | 400K | 23.33 | - | - | - |
| DiT-XL/2 | 400K | 19.47 | - | - | - |
| DiG-S/2 [87] | 400K | 62.06 | 22.81 | 0.39 | 0.56 |
| DiG-S/2 [87] | 400K | 39.50 | 37.21 | 0.51 | 0.63 |
| DiG-S/2 [87] | 400K | 22.90 | 59.87 | 0.60 | 0.64 |
| DiG-S/2 [87] | 400K | 18.53 | 68.53 | 0.63 | 0.64 |
| LiT-S/2 | 100K | 60.91 | 23.51 | 0.399 | 0.583 |
| LiT-B/2 | 100K | 38.39 | 38.59 | 0.516 | 0.627 |
| LiT-L/2 | 100K | 17.80 | 76.26 | 0.639 | 0.636 |
| LiT-XL/2 | 100K | 12.90 | 95.80 | 0.657 | 0.653 |
| LiT-S/2 | 400K | 50.79 | 29.17 | 0.443 | 0.618 |
| LiT-B/2 | 400K | 29.55 | 50.57 | 0.557 | 0.645 |
| LiT-L/2 | 400K | 13.73 | 90.53 | 0.656 | 0.650 |
| LiT-XL/2 | 400K | 10.67 | 110.95 | 0.668 | 0.667 |
| Class-Conditional ImageNet 256256 | ||||
| Model | FID | IS | Precision | Recall |
| BigGAN-deep [2] | 6.95 | 171.4 | 0.87 | 0.28 |
| StyleGAN-XL [66] | 2.30 | 265.12 | 0.78 | 0.53 |
| ADM [14] | 10.94 | 100.98 | 0.69 | 0.63 |
| ADM-G | 4.59 | 186.70 | 0.82 | 0.52 |
| ADM-G, ADM-U | 3.94 | 215.84 | 0.83 | 0.53 |
| CDM [30] | 4.88 | 158.71 | - | - |
| RIN [36] | 3.42 | 182.0 | - | - |
| LDM-4-G (cfg=1.25) [62] | 3.95 | 178.22 | 0.81 | 0.55 |
| LDM-4-G (cfg=1.50) | 3.60 | 247.67 | 0.87 | 0.48 |
| Simple Diffusion (U-Net) [32] | 3.76 | 171.6 | - | - |
| Mask-GIT [5] | 6.18 | 182.1 | - | - |
| Simple Diffusion (U-ViT, L) | 2.77 | 211.8 | - | - |
| DiT-XL/2 [53] | 9.62 | 121.50 | 0.67 | 0.67 |
| DiT-XL/2-G (cfg=1.25) | 3.22 | 201.77 | 0.76 | 0.62 |
| DiT-XL/2-G (cfg=1.50) | 2.27 | 278.24 | 0.83 | 0.57 |
| SiT-XL [49] (cfg=1.50) | 2.06 | 277.50 | 0.83 | 0.59 |
| DiM-L [69] | 2.64 | - | - | - |
| DiM-H [69] | 2.40 | - | - | - |
| DiffuSSM-XL-G [77] | 2.28 | 259.13 | 0.86 | 0.56 |
| LiT-XL/2 | 10.24 | 114.79 | 0.666 | 0.674 |
| LiT-XL/2-G (cfg=1.25) | 3.60 | 191.06 | 0.758 | 0.623 |
| LiT-XL/2-G (cfg=1.50) | 2.32 | 265.20 | 0.824 | 0.574 |
As demonstrated in IDDPM [52], a suitable variance can contribute to the initial steps of the diffusion process, thereby improving the log-likelihood. Accordingly, the authors set the variance learnable. Based on the finding, we propose to use the variance from the teacher diffusion model as a supervision for the student, but in a moderate way. The MSE between the reverse process variances of teacher and student can be calculated as:
where . With Eq. 5 and Eq. 6, we introduce a hybrid diffusion knowledge distillation objective, involving both the predicted noise and the variance:
where , are hyper-parameters. We thoroughly study some representative settings of our hybrid distillation objective as presented in Tab. 4, and identifies two key findings: (1) Compared to the baseline (distilling noise only), the proposed hybrid distillation contributes positively (50.79 vs. 51.13, measured by FID), but the variance distillation should be moderate. (e.g., set to 0.05). (2) Distilling only the variance is counterproductive for optimization, probably as the denoising capability is the major focus for diffusion.
Closing remarks.
So far, we have finalized the roadmap. LiT follows the macro/micro-level design of DiT but employs efficient linear attention. Using our cost-effective training strategies, LiT-S/2 significantly improves the FID. Next, we validate it on larger variants (e.g., B/L/XL) and challenging tasks (e.g., text-to-image generation).
5 Experiments
5.1 Class-Conditional Image Generation
256256 ImageNet.
We first evaluate LiT on the 256256 ImageNet [13] benchmark for conditional image generation. The LiT-S/2, B/2, L/2, XL/2 configurations are consistent with DiT, except that the linear attention heads are set to 2/3/4/4, following the guideline 2 in Sec. 4.2. DWC kernel size is set to 5 for all model variants. Detailed configuration are provided in Sec. A of the appendix. We train 400K steps with a batch size of 256. For LiT-XL/2, we extend training to 1.4M steps (20 of DiT’s [53] 7M steps). We used the AdamW [47] optimizer with a learning rate of and weight decay set to 0. Following Sec. 4.3, we initialize LiT’s parameters using the pre-trained DiT. , in hybrid knowledge distillation, are set to 0.5 and 0.05. Evaluation metrics include FID-50K [27], Inception Score [65] and Precision/Recall [40].
Tab. 5 compares the results of different model variants of LiT and DiT. Notably, LiT trained for only 100K steps already outperforms the DiT trained for 400K steps across various evaluation metrics and different model variants. With extra training up to 400K steps, the model’s performance continues to improve. In Tab. 6, we compare LiT-XL/2 trained for 1.4M steps with high-performance baselines. Despite having only of the training steps of DiT-XL/2, LiT still competes on par with DiT (2.32 vs. 2.27). Besides, LiT competes favorably with several U-Net [63] based baselines. The results demonstrate that the linear attention, when combined with suitable optimization strategies, can be reliably used in image generation applications.
| Class-Conditional ImageNet 512512 | ||||
| Model | FID | IS | Precision | Recall |
| BigGAN-deep [2] | 8.43 | 177.90 | 0.88 | 0.29 |
| StyleGAN-XL [66] | 2.41 | 267.75 | 0.77 | 0.52 |
| ADM [14] | 23.24 | 58.06 | 0.73 | 0.60 |
| ADM-U | 9.96 | 121.78 | 0.75 | 0.64 |
| ADM-G | 7.72 | 172.71 | 0.87 | 0.42 |
| ADM-G, ADM-U | 3.85 | 221.72 | 0.84 | 0.53 |
| Simple Diffusion (U-Net) [32] | 4.28 | 171.0 | - | - |
| Mask-GIT [5] | 7.32 | 156.0 | - | - |
| Simple Diffusion (U-ViT, L) | 4.53 | 205.3 | - | - |
| DiT-XL/2 [53] | 12.03 | 105.25 | 0.75 | 0.64 |
| DiT-XL/2-G (cfg=1.25) | 4.64 | 174.77 | 0.81 | 0.57 |
| DiT-XL/2-G (cfg=1.50) | 3.04 | 240.82 | 0.84 | 0.54 |
| SiT-XL [49] (cfg=1.50) | 2.62 | 252.21 | 0.84 | 0.57 |
| LiT-XL/2 | 14.00 | 92.84 | 0.76 | 0.62 |
| LiT-XL/2-G (cfg=1.50) | 3.69 | 207.97 | 0.85 | 0.53 |
512512 ImageNet.
We further validate LiT-XL/2 on the 512512 ImageNet benchmark. Using the pre-trained DiT-XL/2 [53] as the teacher model, we initialize LiT-XL/2 with its weights. For knowledge distillation, we set and to 1.0 and 0.05, respectively, and train LiT-XL/2 for only 700K steps ( of DiT’s [53] 3M steps). Notably, unlike DiT, which used a batch size of 256, we adopt a smaller batch size of 128, requiring twice the training steps to complete 1 epoch (i.e., processing the entire training dataset once). Aside from this, all other hyper-parameter settings were consistent with the 256256 experiments. The results are reported in Tab. 7.
Despite being trained for only 700K steps, which is equivalent in training epochs to 350K steps with a batch size of 256 on the same dataset, LiT, using pure linear attention, achieves an impressive FID of 3.69, comparable to DiT trained for 3M steps, reducing the training steps by . Additionally, LiT outperforms several strong baselines. These results demonstrate the effectiveness of our proposed cost-efficient training strategy on higher-resolution datasets.
5.2 Text-to-Image Generation
512px resolution.
Text-to-image generation is important in generative models for commercial use. Following PixArt- [8], we add cross-attention module to our LiT-XL/2 to support text embedding input. We set the linear attention heads to 2 and the DWC kernel size to 5. Following PixArt- [7], we use pre-trained SDXL [56] VAE Encoder and T5 [10] Encoder (i.e., Flan-T5-XXL) to extract image and text features, respectively. We use PixArt- as the teacher to supervise the training of LiT, with , set to 1.0, 0.05. As in Sec. 4.3, we inherit the weights from PixArt-, except for the parameters in the self-attention. Subsequently, we train LiT on an internal dataset with a learning rate of for only 45400 steps, which is significantly fewer than the multi-stage training used for PixArt-.
Fig. 7, Fig. 11, and Fig. 12 (illustrated in Sec. E of the appendix) depict the 512px images generated by LiT. Despite using linear attention in each block, along with our cost-effective training strategy, LiT can still produce exceptional, photorealistic images. The results show the generality and commercial value of our approach for cost-effective training of efficient text-to-image frameworks.
1K resolution.
We then scale the resolution to the more challenging 1K setting. Model of 512px is selected as the starting point for training, with positional embeddings (PE) Interpolation trick applied. Token length is increased from to . We use the PixArt- [7] model, designed for 1K resolution, as the teacher model. We prepare two internal datasets: one follows the PixArt- setup, and the other follows the PixArt- setup, with a caption ratio of long (sharegpt caption) and short (original captions). The training consists of multiple stages: stage 1 trains on the first dataset, and stages 2, 3, and 4 train on the second dataset. We use 32 NVIDIA V100 GPUs, with batch sizes of 64, 64, 256, and 256 for the four stages. In the final stage, no knowledge distillation method is used.
Fig. 1 shows the sampled 1K resolution images generated by LiT. Despite replacing all self-attention with the cheap linear attention, LiT is still able to generate photo-realistic images at high resolution.
5.3 On-Device Deployment
In this section, we deploy the LiT model for 1K resolution images to an edge device to validate its on-device capability. Specifically, we deploy the model trained after stage 3 of the 1K resolution experiments from Sec. 5.2 to a laptop powered by the Windows 11 operating system. Considering the GPU memory limitations of the laptop, we quantize the text encoder to 8 bits while maintaining fp16 precision during the linear attention computation. Fig. 6 shows our deployment results. The pre-trained LiT can quickly generate photo-realistic 1K resolution images in an offline setting (without network connection). These results confirm the successful implementation of LiT as an on-device diffusion model, advancing high-resolution text-to-image tasks on edge devices.
6 Conclusion and Limitation
This paper explores using linear attention to implement diffusion Transformer for text-to-image task. We reexamine attention head number, weight inheritance, and knowledge distillation. Our model can handle class-conditional image generation, as well as synthesize photorealistic images based on instructions. Ultra-high resolution scenarios (e.g., 4K resolution) for image generation remains a limitation, which we leave in future work.
Supplementary Material
Appendix A Model Configuration
We report the configurations of LiT variants in Tab. 8, which basically follow the hyper-parameters of DiT [53], except for using few heads. For class-conditional image generation, we set 2/3/4/4 heads for linear attention used in LiT. For LiT-XL/2, used for text-to-image tasks, we use 2 heads for linear attention.
| Model | Layers | Hidden size | Heads | Patch size |
| LiT-S | 12 | 384 | 2 | 2 |
| LiT-B | 12 | 768 | 3 | 2 |
| LiT-L | 24 | 1024 | 4 | 2 |
| LiT-XL | 28 | 1152 | 4 | 2 |
| LiT-XL⋄ | 28 | 1152 | 2 | 2 |
Appendix B Latency Analysis in Diffusion Transformer
We conduct a component-wise latency analysis of the Diffusion Transformer, with results shown in Fig. 8. The analysis was performed using the DiT-B/4 [53] model on an NVIDIA A100 GPU. The results indicate that the self-attention module accounts for 42.6% of the total latency of a DiT block. We attribute the observed considerable latency proportion to the quadratic computational complexity of the self-attention.
Appendix C Detailed Latency and Theoretical GMACs
We use one NVIDIA V100 GPU to evaluate the latency and theoretical GMACs of the DiT-S/B models with different numbers of heads. The task was to generate 256256 resolution images with a batch size of 8, following the experimental setup of Fig. 5 (except for the GPU type), and the results are shown in Fig. 10.
We observe that both the small and base models on the V100 GPU exhibit a phenomenon similar to that on the A100 GPU: as theoretical GMACs increased, practical latency does not follow the same trend and even descend (in the B/2 model). This finding supports the generalizability of the free lunch effect of linear attention.
Appendix D Detailed Results on Class-Conditional Image Generation
Detailed results on Sec. 4.
We provide detailed results for Tab. 1, Tab. 2, Tab. 3 and Tab. 4 in the main text in Tab. 9, Tab. 10, Tab. 11, and Tab. 12, respectively. In each table, we report results involving FID-50K [27] (without classifier-free guidance), Inception Score (IS) [65] and Precision/Recall [40]. IS, and Precision/Recall show results similar to FID-50K. As a result, the conclusions drawn in Sec. 4 of the main paper apply not only to metrics evaluating the distance between generated images and real images (e.g., FID-50K) but also to metrics reflecting the quality of the generated images themselves (e.g., IS).
Appendix E More Results on Text-to-Image Generation
We provide more text-to-image results in Fig. 11 and Fig. 12. As shown, LiT can accurately generate 512px photo-realistic images in various styles, themes, and content, whether the human instructions are simple or complicated. These results demonstrate that LiT effectively learns useful knowledge from the teacher model while maintaining exceptional computational efficiency, validating the effectiveness of our proposed cost-effective training strategy.
Appendix F Code in PyTorch
PyTorch-like code of the linear attention used in our LiT for text-to-image generation is provided in Alg. 1. As shown, when training LiT for text-to-image generation, we use fp16 precision (except for the main part of linear attention, where fp32 is used) to prevent the loss from becoming NaN, basically following PixArt- [7]. We set the kernel size of the depthwise convolution to 5.
Appendix G Full Related Work
Linear attention.
As a computationally efficient alternative to self-attention, linear attention [38] reduces computational complexity from quadratic to linear and has been proven effective in both visual understanding domain [3, 22, 23, 24, 1] and language domain [59, 78, 61, 60, 6]. EfficientViT [3] introduces a multi-scale linear attention with hardware-efficient operations to obtain a general vision backbone. Flatten Transformer [23] introduces focused linear attention to address the deficiencies in focus ability and feature diversity of linear attention, incorporating a focused function and depthwise convolution (DWC). SLAB [22] simplifies focused linear attention by retaining only the DWC component and introduces a progressive re-parameterized batch normalization to adapt offline batch normalization [35] for achieving low inference latency. These studies have been validated on visual understanding. Meanwhile, our work refines a linear attention module tailored for image generation tasks and identifies the free lunch of using few heads.
Efficient diffusion Transformer for image generation.
Limited by the quadratic computational complexity of self-attention, recent studies focus on developing efficient Transformer-style architectures for diffusion models. For example, DiM [69], ZigMa [34], and DiMSUM [55] explore implementing Mamba-based [21, 11] DiT-style [53] models. Diffusion-RMKV [17] studies RWKV-style [54] models for diffusion. Mediator [58] introduces an attention mediator to obtain an efficient diffusion Transformer with linear complexity. DiG [87] replaces the self-attention in DiT with gated linear attention to speed up training. LinFusion [46] and Sana [75] apply linear attention to U-Net-based [63] and Transformer-based [71] diffusion models, respectively, and train these models from scratch to generate high-quality images based on user instructions. Other studies [43, 42] explore efficient diffusion models through perspectives of low-bit quantization [25], feature map reusing [50, 41], and lightweight architecture design [86]. Differently, our work not only refines the design of linear attention but also introduces cost-effective training strategies, including weight inheritance and a novel hybrid diffusion distillation approach.
Advanced training method for diffusion models.
Some studies explores improved training strategies to enhance the optimization of diffusion models or achieve more efficient models. For example, CAN [4] introduces a condition-aware weight generation module to the diffusion Transformer, and demonstrate the technique can be further equipped with EffcientViT [3] to achieve both effectiveness and efficiency. REPA [82] proposes aligning the intermediate features of the denoising model with those extracted by a pre-trained visual encoder during the training of diffusion models. Some studies [64, 81, 80] explore distillation techniques to reduce the sampling steps of the diffusion model. Unlike the goal of reducing sampling steps, our proposed hybrid knowledge distillation focuses on an architectural perspective, exploring how a lightweight student model can learn from a computationally intensive teacher model.
| DiT | Attention | Resolution | Batch Size | Training Steps | FID-50K () | IS () | Precision () | Recall () |
| S/2 | softmax | 256 | 256 | 400K | 68.40 | - | ||
| S/2 | ReLU linear | 256 | 256 | 400K | 88.46 | 15.11 | 0.29 | 0.45 |
| S/2 | Simplified linear (ReLU) | 256 | 256 | 400K | 63.66 | 22.16 | 0.38 | 0.58 |
| S/2 | focused linear (ReLU) | 256 | 256 | 400K | 63.05 | 22.49 | 0.39 | 0.58 |
| S/2 | focused linear (GELU) | 256 | 256 | 400K | 70.83 | 19.41 | 0.36 | 0.54 |
| B/2 | softmax | 256 | 256 | 400K | 43.47 | - | ||
| B/2 | ReLU linear | 256 | 256 | 400K | 56.92 | 25.80 | 0.42 | 0.59 |
| B/2 | Simplified linear (ReLU) | 256 | 256 | 400K | 42.11 | 34.60 | 0.48 | 0.63 |
| B/2 | focused linear (ReLU) | 256 | 256 | 400K | 40.58 | 35.98 | 0.50 | 0.63 |
| B/2 | focused linear (GELU) | 256 | 256 | 400K | 58.86 | 24.23 | 0.42 | 0.57 |
| DiT | Head | Resolution | Batch Size | Training Steps | FID-50K () | IS () | Precision () | Recall () |
| S/2 | 1 | 256 | 256 | 400K | 64.42 | 21.54 | 0.380 | 0.574 |
| S/2 | 2 | 256 | 256 | 400K | 63.24 | 22.07 | 0.385 | 0.570 |
| S/2 | 3 | 256 | 256 | 400K | 63.21 | 22.08 | 0.386 | 0.583 |
| S/2 | 6 | 256 | 256 | 400K | 63.66 | 22.16 | 0.383 | 0.580 |
| S/2 | 48 | 256 | 256 | 400K | 78.76 | 17.46 | 0.322 | 0.482 |
| S/2 | 96 | 256 | 256 | 400K | 116.00 | 11.49 | 0.224 | 0.261 |
| B/2 | 1 | 256 | 256 | 400K | 41.77 | 34.78 | 0.487 | 0.631 |
| B/2 | 2 | 256 | 256 | 400K | 41.39 | 35.59 | 0.494 | 0.631 |
| B/2 | 3 | 256 | 256 | 400K | 40.86 | 35.79 | 0.497 | 0.629 |
| B/2 | 12 | 256 | 256 | 400K | 42.11 | 34.60 | 0.484 | 0.631 |
| B/2 | 96 | 256 | 256 | 400K | 68.30 | 20.45 | 0.375 | 0.531 |
| B/2 | 192 | 256 | 256 | 400K | 112.39 | 12.07 | 0.240 | 0.282 |
| L/2 | 1 | 256 | 256 | 400K | 24.46 | 57.36 | 0.600 | 0.637 |
| L/2 | 2 | 256 | 256 | 400K | 24.37 | 57.02 | 0.599 | 0.622 |
| L/2 | 4 | 256 | 256 | 400K | 24.04 | 59.02 | 0.597 | 0.636 |
| L/2 | 16 | 256 | 256 | 400K | 25.25 | 54.67 | 0.587 | 0.632 |
| XL/2 | 1 | 256 | 256 | 400K | 21.13 | 65.06 | 0.619 | 0.632 |
| XL/2 | 2 | 256 | 256 | 400K | 20.66 | 65.39 | 0.624 | 0.636 |
| XL/2 | 4 | 256 | 256 | 400K | 20.82 | 65.52 | 0.619 | 0.632 |
| XL/2 | 16 | 256 | 256 | 400K | 21.69 | 63.06 | 0.617 | 0.628 |
| Load | Iterations | FFN | Modulation Layer | Attention | FID-50K () | IS () | Precision () | Recall () |
| model | 400K | ✓ | ✓ | ✗ | 56.07 | 25.62 | 0.418 | 0.608 |
| ema | 400K | ✓ | ✓ | ✗ | 56.07 | 25.61 | 0.416 | 0.601 |
| model | 200K | ✓ | ✓ | ✗ | 57.84 | 24.72 | 0.408 | 0.600 |
| model | 300K | ✓ | ✓ | ✗ | 56.95 | 25.04 | 0.414 | 0.608 |
| model | 400K | ✓ | ✓ | ✗ | 56.07 | 25.62 | 0.418 | 0.608 |
| model | 600K | ✓ | ✓ | ✗ | 54.80 | 26.65 | 0.424 | 0.613 |
| model | 800K | ✓ | ✓ | ✗ | 53.83 | 27.16 | 0.425 | 0.614 |
| model | 600K | ✓ | ✓ | Q, K, V | 55.29 | 26.09 | 0.419 | 0.619 |
| model | 600K | ✓ | ✓ | K, V | 55.07 | 26.38 | 0.422 | 0.609 |
| model | 600K | ✓ | ✓ | V | 54.93 | 26.44 | 0.427 | 0.612 |
| model | 600K | ✓ | ✓ | Q | 54.82 | 26.72 | 0.423 | 0.605 |
| model | 600K | ✓ | ✓ | O | 54.84 | 26.33 | 0.425 | 0.607 |
| Model Size | Iterations | Teacher | Training Steps | FID-50K () | IS () | Precision () | Recall () | ||
| S/2 | 800K | DiT-S/2 | 0.1 | 0.0 | 400K | 55.11 | 26.28 | 0.419 | 0.614 |
| S/2 | 800K | DiT-XL/2 | 0.0 | 0.0 | 400K | 53.83 | 27.16 | 0.425 | 0.614 |
| S/2 | 800K | DiT-XL/2 | 0.1 | 0.0 | 400K | 53.05 | 27.43 | 0.431 | 0.609 |
| S/2 | 800K | DiT-XL/2 | 0.05 | 0.0 | 400K | 53.41 | 27.26 | 0.427 | 0.610 |
| S/2 | 800K | DiT-XL/2 | 0.5 | 0.0 | 400K | 51.13 | 28.89 | 0.438 | 0.616 |
| S/2 | 800K | DiT-XL/2 | 0.1 | 0.05 | 400K | 52.76 | 27.70 | 0.431 | 0.620 |
| S/2 | 800K | DiT-XL/2 | 0.0 | 0.05 | 400K | 53.49 | 27.26 | 0.429 | 0.609 |
| S/2 | 800K | DiT-XL/2 | 0.05 | 0.05 | 400K | 53.14 | 27.46 | 0.431 | 0.609 |
| S/2 | 800K | DiT-XL/2 | 0.5 | 0.05 | 400K | 50.79 | 29.17 | 0.443 | 0.618 |