### v2

# LiT: Delving into a Simple Linear Diffusion Transformer for Image Generation

Abstract
In this paper, we investigate how to convert a pre-trained Diffusion Transformer (DiT) into a linear DiT, as its simplicity, parallelism, and efficiency for image generation. Through detailed exploration, we offer a suite of ready-to-use solutions, ranging from linear attention design to optimization strategies. Our core contributions include 5 practical guidelines: 1) Applying depth-wise convolution within simple linear attention is sufficient for image generation. 2) Using fewer heads in linear attention provides a free-lunch performance boost without increasing latency. 3) Inheriting weights from a fully converged, pre-trained DiT. 4) Loading all parameters except those related to linear attention. 5) Hybrid knowledge distillation: using a pre-trained teacher DiT to help the training of the student linear DiT, supervising not only the predicted noise but also the variance of the reverse diffusion process. These guidelines lead to our proposed Linear Diffusion Transformer (LiT), which serves as a safe and efficient alternative baseline for DiT with pure linear attention. In class-conditional 256256 and 512512 ImageNet generation, LiT can be quickly adapted from DiT using only and of DiT’s training steps, respectively, while achieving comparable performance. LiT also rivals methods based on Mamba or Gated Linear Attention. Moreover, the same guidelines generalize to text-to-image generation: LiT can be swiftly converted from PixArt- to generate high-quality images, maintaining comparable GenEval scores.
1 Introduction
Diffusion models [30, 68] based on Transformers [7, 16, 8] show strong commercial potential in image generation but face latency and GPU memory challenges in high-resolution tasks. Their main self-attention modules exhibit quadratic complexity with respect to sequence length, making deployment on resource-limited edge devices less feasible, as illustrated in Fig. 4 and Fig. 1 in Appendix Sec. B, highlighting the need for sub-quadratic attention modules.
Linear attention can significantly accelerate image generation compared to self-attention at the same resolution. For example, at 2048px, it is nearly 9 faster (Fig. 4-(a), (b)). Additionally, it substantially reduces GPU memory usage for high-resolution image generation. Replacing self-attention with linear attention in DiT-S/2 reduces memory consumption from 14GB to 4GB at 2048px. In fact, linear attention has been extensively validated in visual recognition tasks (e.g., image classification and segmentation) [3, 24, 1]. In these methods, the linear attention is expected to generate sharp distributions that effectively capture semantic information [3, 24]. However, in the context of denoising diffusion models, linear attention may require enhanced locality to better model the distribution of random noise. Consequently, exploring the use of linear attention in image generation remains a promising research direction.
Several efforts have explored the use of efficient attention for image generation. For instance, DiG [88] and DiM [71] reference Gated Linear Attention (GLA) [80] and State Space Models (SSM) [21, 22] to implement diffusion models, respectively. SANA [78] takes a deep compression autoencoder [9], and replace the original MLP-FFN with Mix-FFN [3]. Attention Mediators [59] applies a set of additional tokens in linear attention. CLEAR [46] adopts a circular window within the attention.
In this paper, we systematically investigate the problem of converting a pre-trained DiT to a linear DiT for image generation while minimizing performance loss. We decompose this problem into two parts: first, what type of linear attention architecture is suitable for image generation? Second, what training strategies are effective for optimizing a linear DiT? To address these two questions, we conducted extensive experiments on Linear DiT from both architecture design and training strategy perspectives. Based on the results, our findings can be summarized into five practical guidelines for Linear DiT design and optimization: 1) simply adding a depth-wise convolution in linear attention is sufficient for image generation; 2) a free lunch: few (e.g., 2) heads in linear attention improves performance without increasing latency; 3) initialize the weights of the linear DiT from a well-trained DiT; 4) load all weights except for the linear attention to achieve cost-effective training; 5) hybrid distillation objective—distilling not only the predicted noise but also variances of the reverse diffusion process—to facilitate the student linear DiT. Following such guidelines (Fig. 3), the result diffusion model, dubbed Linear Diffusion Transformer (LiT), can adopt pure simple attention design, achieving linear computational complexity while retaining the capability of full attention. Meanwhile, LiT is trained in a highly cost-effective manner as it is derived from a pre-trained DiT, leveraging existing knowledge and reducing the need for extensive training from scratch.
LiT is an affirmative answer to how to efficiently and safely obtain a linear DiT from a pretrained DiT. To validate its effectiveness, we conduct experiments on representative diffusion Transformers, i.e., DiT (for ImageNet [14] class-conditional image generation) and PixArt- [7] (for text-to-image generation). In ImageNet [14] 256256 benchmark, LiT-S/B/L (patch size of 2) with only 100K training steps outperforms corresponding 400K steps pre-trained DiT-S/B/L in terms of Fréchet Inception Distance (FID) [28]. For large-scale model on the ImageNet 256256 and 512512 benchmark, LiT-XL/2 competes comparably to DiT-XL/2 (2.32 vs. 2.27, 3.69 vs. 3.04 in FID) with only and of training steps (respectively for 256256 and 512512 settings). In text-to-image generation, converted from PixArt-, LiT-0.6B shares the same macro/micro-level architecture, ensuring alignment with the PixArt- framework while elegantly replacing all self-attention with efficient linear attention. LiT-0.6B can stably generate highly photorealistic images (Fig. 2). To verify the effectiveness of linear attention, we deploy LiT-0.6B offline on a Windows 11 laptop (Fig. 11), enabling 1K-resolution image synthesis based on human instructions.
2 Preliminaries
Diffusion models.
Suppose we have a clean image , and progressively add noise at each step to obtain . The forward (diffusion) process is defined as: , where and . Diffusion models [30, 68, 49, 77] learn a network to reverse the diffusion process: , where and denote the mean and variance of the reverse process. In DDPM [30], a simple training objective is demonstrated effective in training diffusion models: , where represents noise of the step, expected to be approximated by the output of a denoising network . can be realized using Transformers [54, 50, 8, 7] with time-intensive self-attention, calling for cheap alternatives.
| DiT | Attention | FID-50K () | IS () |
| S/2 | Softmax | 68.40 | - |
| S/2 | ReLU Linear Baseline | 88.46 | 15.11 |
| S/2 | + Depth-wise Conv. (ReLU) | 63.66 | 22.16 |
| S/2 | + Focused Linear (ReLU) | 63.05 | 22.49 |
| S/2 | + Focused Linear (GELU) | 70.83 | 19.41 |
| B/2 | softmax | 43.47 | - |
| B/2 | ReLU Linear Baseline | 56.92 | 25.80 |
| B/2 | + Depth-wise Conv. (ReLU) | 42.11 | 34.60 |
| B/2 | + Focused Linear (ReLU) | 40.58 | 35.98 |
| B/2 | + Focused Linear (GELU) | 58.86 | 24.23 |
Linear attention.
For a standard softmax attention with heads, sequence length , and hidden dimension , denoting as query, key and value, its output can be expressed as:
where is a similarity measurement function. In linear attention, the aforementioned function is modified into a simplified form with a nonlinear kernel , i.e., . Thus, Eq. 1 can be rewritten as Eq. 2 and simplified to Eq. 3 using the associative property of multiplication.
The computational complexities of softmax attention and linear attention are and , respectively.
Despite its efficiency, the suitability of linear attention for image generation remains under-explored. Most existing works, such as SANA [78], Mediators [59], and CLEAR [46], introduce carefully designed modifications to DiT, compromising its simplicity. Motivated by this, we explore how to safely and efficiently convert a pretrained DiT into a linear DiT without relying on complex architectures or sacrificing performance.
3 Exploring LiT: A Roadmap
We explore the architectural design and optimization strategies of our LiT based on DiT [54] on class-conditional ImageNet [14] 256256 benchmark, as it is relatively simple, thereby highlighting the ablation effect of attention itself. We train models for 400K iterations using a batch size of 256, and report the FID-50K [28], Inception Score [66] and Precision/Recall [40] metrics. The radar chart of the experimental results is shown in Fig. 1. Detailed results can be found in Sec. E of the Appendix.
3.1 Architectural Guidelines for Linear DiT
| (a) Latency | (b) GPU Memory |
| DiT | Head | FID-50K () | IS () | Prec. () | Rec. () |
| S/2 | 1 | 64.42 | 21.54 | 0.380 | 0.574 |
| S/2 | 2 | 63.24 | 22.07 | 0.385 | 0.570 |
| S/2 | 3 | 63.21 | 22.08 | 0.386 | 0.583 |
| S/2 | 6 | 63.66 | 22.16 | 0.383 | 0.580 |
| S/2 | 48 | 78.76 | 17.46 | 0.322 | 0.482 |
| S/2 | 96 | 116.00 | 11.49 | 0.224 | 0.261 |
| B/2 | 1 | 41.77 | 34.78 | 0.487 | 0.631 |
| B/2 | 2 | 41.39 | 35.59 | 0.494 | 0.631 |
| B/2 | 3 | 40.86 | 35.79 | 0.497 | 0.629 |
| B/2 | 12 | 42.11 | 34.60 | 0.484 | 0.631 |
| B/2 | 96 | 68.30 | 20.45 | 0.375 | 0.531 |
| B/2 | 192 | 112.39 | 12.07 | 0.240 | 0.282 |
| L/2 | 1 | 24.46 | 57.36 | 0.600 | 0.637 |
| L/2 | 2 | 24.37 | 57.02 | 0.599 | 0.622 |
| L/2 | 4 | 24.04 | 59.02 | 0.597 | 0.636 |
| L/2 | 16 | 25.25 | 54.67 | 0.587 | 0.632 |
| XL/2 | 1 | 21.13 | 65.06 | 0.619 | 0.632 |
| XL/2 | 2 | 20.66 | 65.39 | 0.624 | 0.636 |
| XL/2 | 4 | 20.82 | 65.52 | 0.619 | 0.632 |
| XL/2 | 16 | 21.69 | 63.06 | 0.617 | 0.628 |
Guideline 1: Simply adding a depth-wise convolution in linear attention is sufficient for DiT-based image generation.
In high-level visual recognition tasks (e.g., image classification, instance segmentation, etc.), linear attention has been reported to produce a “smooth” attention distribution compared to the “sharp” distribution of softmax attention [3, 24]. To address the issue, Flatten Transformer [24] introduces a focused function applied after the ReLU activation as a remedy. Meanwhile, linear attention has been shown to exhibit limited feature diversity due to its low-rank attention matrix. Accordingly, only a single depth-wise convolution layer, applied on the value matrix, can raise the upper bound of the equivalent rank, effectively mitigating the low-rank issue [24]. Based on these observations, it is reasonable to validate similar conclusions in image generation. Compared to the full attention baseline, LiT-S/2 and B/2 with ReLU linear attention exhibit an unacceptable performance drop (Tab. 1), with FID increasing by approximately 20/13 for the S/2 and B/2 models, respectively. We then explore the following approaches: (1) adding a depth-wise convolution (DWC) [34, 10] on the value matrix111In this setup, we optimize the linear attention implementation (e.g., by replacing einsum with the “@” operator for matrix operations and removing the softplus function). We also test a version without these optimizations for comparison, which yielded results of 63.13 (S/2) and 42.13 (B/2), showing only slight differences from the results in Tab. 1: 63.66 (S/2) and 42.11 (B/2). Therefore, we retain the optimizations.; (2) adding focused function [24]; (3) replacing the ReLU kernel with GELU [27] function, commonly used in Transformers [73]. Each of these choices maintains linear complexity, preserving LiT’s advantage in computational efficiency. We use a relatively large convolution kernel (i.e., 5) to ensure a sufficiently large receptive field when predicting noise.
As shown in Tab. 1, incorporating DWC alone enhances generation quality beyond softmax attention, achieving FID scores of 63.66 and 42.11 for S/2 and B/2, respectively. We attribute this improvement to the model’s tendency to rely on information from neighboring pixels when predicting noise for a given pixel, necessitating the assistance of convolution. Meanwhile, we identify that the focused function has limited effectiveness, which we attribute to its design motivation to help linear attention focus on specific regions. This feature may suit classification models, but may not be necessary for noise prediction.
Remark 1. We note that our findings on linear attention in visual generation align with those of the SLAB [23] in visual recognition, where simple DWC leads to improvement. We attribute this to linear attention’s lack of locality modeling capacity compared to full attention, which DWC serves as a remedy. We will adopt this design afterwards.
| Load | Iter. | FFN | Modu. | Attention | FID-50K () |
| model | 400K | ✓ | ✓ | ✗ | 56.07 |
| ema | 400K | ✓ | ✓ | ✗ | 56.07 |
| model | 200K | ✓ | ✓ | ✗ | 57.84 |
| model | 300K | ✓ | ✓ | ✗ | 56.95 |
| model | 400K | ✓ | ✓ | ✗ | 56.07 |
| model | 600K | ✓ | ✓ | ✗ | 54.80 |
| model | 800K | ✓ | ✓ | ✗ | 53.83 |
| model | 600K | ✓ | ✓ | Q, K, V | 55.29 |
| model | 600K | ✓ | ✓ | K, V | 55.07 |
| model | 600K | ✓ | ✓ | V | 54.93 |
| model | 600K | ✓ | ✓ | Q | 54.82 |
| model | 600K | ✓ | ✓ | O | 54.84 |
3.2 Linear Attention with Few Heads
| (a) Small/Base | (b) Large/XLarge |
Guideline 2: Using few heads in the linear attention increases computation but not latency.
In practice, DiT-S/B/L/XL are designed to have 6/12/16/16 heads, respectively. Denoting the number of heads, sequence length, and hidden dimension as , GMACs of multi-head self-attention and linear attention are and , respectively, as follows:
Theoretically, is independent of the number of heads, but in linear attention, has a negative correlation with it. Intuitively, using many heads can loose the computational stress. But on the contrary, we prefer to use few heads, based on the identified free lunch effect of linear attention. We illustrate the effect in Fig. 5, where we visualize the latency and theoretical GMACs of linear attention in DiT-S/B/L/XL models with different head numbers. We use one NVIDIA A100 GPU to generate 256256 resolution images with a batch size of 8, averaging the results over 30 experiments. The results show that decreasing the number of heads leads to a steady increase in theoretical GMACs but not the practical latency, which even shows a decrease. Accordingly, we summarize it as the free lunch effect of linear attention. We also observe this phenomenon on NVIDIA V100 GPU, as reported in Fig. 3 in Appendix Sec. C. As such, we argue that setting a small number of heads enables a high theoretical compute budget, which, according to scaling law [44, 38, 32], allows the model to reach a higher upper bound in generation performance. The argument can be empirically demonstrated in Tab. 2. For different model scales, equipping linear attention with a few heads (e.g., 2, 3, 4) outperforms the default settings in DiT. Instead, using an excessive number of heads (e.g., 96 for S/2 or 192 for B/2) seriously hinders the generation quality.
Besides, we also evaluate the necessity of using many heads in linear attention from a representation similarity lens. Specifically, we visualize the average cosine similarity among the attention maps of different heads in LiT-S/2 (with 6 heads). As depicted in Fig. 4 in Appendix Sec. D, for different denoising steps or layers, the attention across different heads is quite homogeneous, with the average cosine similarity reaching up to above 0.5. The observation suggests that a small number of heads dominate the major information in the linear attention, which positively supports the empirical generative results in Tab. 2.
3.3 Inheriting Weights from DiT
LiT can be converted from a pre-trained DiT, whose weights contain rich knowledge of noise prediction, enabling a cost-effective transfer to LiT. Thus, we load the pre-trained DiT weights into LiT, except for the linear attention. DiT can be pretrained to varying convergence levels. Given the architectural differences between DiT and LiT, whether converged DiT weights serve as optimal initialization for LiT remains an open question. To this end, we conduct ablation experiments to explore how the degree of training on inherited weights affects the model’s generation performance.
Guideline 3: Linear diffusion Transformer should be initialized from a converged DiT.
We first pre-train DiT-S/2 [54] for five types of iterations—200K, 300K, 400K, 600K, and 800K—and in separate experiments, load these pre-trained weights into LiT-S/2, keeping only the linear attention parameters randomly initialized. The initialized LiT-S/2 is then trained for 400K iterations on class-conditional ImageNet, with results shown in Tab. 3. Some interesting findings are identified: (1) DiT’s pre-trained weights, even trained only 200K steps, play a significant role, improving the FID from 63.24 to 57.84. (2) Using exponential moving average (EMA) [58] of pre-trained weights has minimal impact. (3) Longer pre-training of DiT makes them better suited as initialization for LiT, even though the architectures are not fully aligned.
Remark 3. We suppose a possible explanation for this phenomenon is the functional decoupling of different modules in DiTs, as explored in large language models [81, 75]. Basically, although DiT and LiT have different architectures, their shared components (e.g., FFN and adaLN) act quite similarly. As a result, knowledge in these components can be transferred, and extensive pre-training may not hinder the optimization of the non-transfered parts.
| Iter. | Teacher | FID-50K () | IS () | ||
| 800K | DiT-S/2 | 0.1 | 0.0 | 55.11 | 26.28 |
| 800K | DiT-XL/2 | 0.0 | 0.0 | 53.83 | 27.16 |
| 800K | DiT-XL/2 | 0.1 | 0.0 | 53.05 | 27.43 |
| 800K | DiT-XL/2 | 0.05 | 0.0 | 53.41 | 27.26 |
| 800K | DiT-XL/2 | 0.5 | 0.0 | 51.13 | 28.89 |
| 800K | DiT-XL/2 | 0.1 | 0.05 | 52.76 | 27.70 |
| 800K | DiT-XL/2 | 0.0 | 0.05 | 53.49 | 27.26 |
| 800K | DiT-XL/2 | 0.05 | 0.05 | 53.14 | 27.46 |
| 800K | DiT-XL/2 | 0.5 | 0.05 | 50.79 | 29.17 |
Guideline 4: Projection matrices of query, key, value, and output in linear attention should be initialized randomly.
Linear attention shares the query, key, value, and output projection matrices with full attention. Despite differing computation paradigms, these weights can be directly transferred from DiT to LiT without modification. However, as shown in Fig. 6, it remains an open question whether this can accelerate convergence. We use a DiT-S/2 pre-trained for 600K iterations for ablation. We explore five different loading strategies, i.e., (1) , , and projection matrices; (2) , and projection matrices; (3) projection matrix; (4) projection matrix; and (5) the output projection matrix. Evaluation results are presented in Tab. 3. Compared to the baseline without self-attention weight loading, none of the explored strategies improve generative performance. This can be attributed to differences in computational paradigms: linear attention directly computes the product of key and value matrices, unlike self-attention. As a result, self-attention’s key and value weights provide limited benefit to linear attention. As shown in Tab. 3, inheriting these weights leads to greater performance degradation, with FID scores of 55.29 and 55.07.
Remark 4. Inheriting weights from a LiT-like pre-trained model is a favorable approach. We recommend loading all pre-trained parameters except for the linear attention, as it is easy to implement and well-suited for diffusion models based on the Transformer macro architecture. With substantial GPU resources, extensive pre-training can further improve linear LiT’s convergence and optimization, resulting in an efficient but strong model.
| Class-Conditional ImageNet 256256 | |||||
| Model | Training Steps | FID | IS | Prec. | Rec. |
| DiT-S/2 [54] | 400K | 68.40 | - | - | - |
| DiT-B/2 | 400K | 43.47 | - | - | - |
| DiT-L/2 | 400K | 23.33 | - | - | - |
| DiT-XL/2 | 400K | 19.47 | - | - | - |
| DiG-S/2 [88] | 400K | 62.06 | 22.81 | 0.39 | 0.56 |
| DiG-B/2 [88] | 400K | 39.50 | 37.21 | 0.51 | 0.63 |
| DiG-L/2 [88] | 400K | 22.90 | 59.87 | 0.60 | 0.64 |
| DiG-XL/2 [88] | 400K | 18.53 | 68.53 | 0.63 | 0.64 |
| LiT-S/2 | 100K | 60.91 | 23.51 | 0.399 | 0.583 |
| LiT-B/2 | 100K | 38.39 | 38.59 | 0.516 | 0.627 |
| LiT-L/2 | 100K | 17.80 | 76.26 | 0.639 | 0.636 |
| LiT-XL/2 | 100K | 12.90 | 95.80 | 0.657 | 0.653 |
| LiT-S/2 | 400K | 50.79 | 29.17 | 0.443 | 0.618 |
| LiT-B/2 | 400K | 29.55 | 50.57 | 0.557 | 0.645 |
| LiT-L/2 | 400K | 13.73 | 90.53 | 0.656 | 0.650 |
| LiT-XL/2 | 400K | 10.67 | 110.95 | 0.668 | 0.667 |
3.4 Distilling Both Predicted Noise and Variance
Guideline 5: Hybrid distillation is neccessary for student linear diffusion Transformer. We distill not only the predicted noise but also variances of the reverse diffusion process, but in a moderate way.
Knowledge distillation [29] commonly employs a teacher network to help the training of a lightweight student network, which has been validated in ConvNets [20] and ViTs [74, 72, 86]. For diffusion models, distillation typically focuses on reducing the sampling steps of the target model [83, 82, 65]. In contrast, we focus on how a heavy teacher model can aid the learning of an efficient student from an architecture perspective, while maintaining the sampling steps [47].
A pre-trained DiT-S/2/XL/2 serves as the teacher network to distill the student LiT-S/2. We use a DiT-S/2 pre-trained for 800K steps for weight initialization of LiT, following Guideline 3&4. Denote as the noised input latent in timestep , where , , are height, width and the number of channels of the input latent. denote network parameters of the teacher diffusion Transformer and the student LiT network , respectively. As denoising diffusion models, the output of DiT and LiT include both noise predictions and variances of the reverse diffusion process . DDPM [30] fixes the variance and optimizes model parameters using . Naturally, a straightforward idea is to align the student network’s noise prediction capabilities with those of the teacher through a mean squared error (MSE) objective:
where is the normalization factor. The impact of distilling the predicted noise is reported in Tab. 4. It shows that using a stronger teacher model (e.g., DiT-XL/2) improves the student model’s generation capbility (51.13 vs. 53.83, measured by FID). Moreover, increasing the weight of the distillation objective from 0.05 to 0.5 consistently boosts performance, suggesting that LiT can benefit from the supervision of teacher DiT.
| Class-Conditional ImageNet 256256 | ||||
| Model | FID | IS | Precision | Recall |
| BigGAN-deep [2] | 6.95 | 171.4 | 0.87 | 0.28 |
| StyleGAN-XL [67] | 2.30 | 265.12 | 0.78 | 0.53 |
| ADM [15] | 10.94 | 100.98 | 0.69 | 0.63 |
| ADM-G | 4.59 | 186.70 | 0.82 | 0.52 |
| ADM-G, ADM-U | 3.94 | 215.84 | 0.83 | 0.53 |
| CDM [31] | 4.88 | 158.71 | - | - |
| RIN [37] | 3.42 | 182.0 | - | - |
| LDM-4-G (cfg=1.25) [63] | 3.95 | 178.22 | 0.81 | 0.55 |
| LDM-4-G (cfg=1.50) | 3.60 | 247.67 | 0.87 | 0.48 |
| Simple Diffusion (U-Net) [33] | 3.76 | 171.6 | - | - |
| Mask-GIT [5] | 6.18 | 182.1 | - | - |
| Simple Diffusion (U-ViT, L) | 2.77 | 211.8 | - | - |
| DiT-XL/2 [54] | 9.62 | 121.50 | 0.67 | 0.67 |
| DiT-XL/2-G (cfg=1.25) | 3.22 | 201.77 | 0.76 | 0.62 |
| DiT-XL/2-G (cfg=1.50) | 2.27 | 278.24 | 0.83 | 0.57 |
| SiT-XL [50] (cfg=1.50) | 2.06 | 277.50 | 0.83 | 0.59 |
| DiM-L [71] | 2.64 | - | - | - |
| DiM-H [71] | 2.40 | - | - | - |
| DiffuSSM-XL-G [79] | 2.28 | 259.13 | 0.86 | 0.56 |
| LiT-XL/2 | 10.24 | 114.79 | 0.666 | 0.674 |
| LiT-XL/2-G (cfg=1.25) | 3.60 | 191.06 | 0.758 | 0.623 |
| LiT-XL/2-G (cfg=1.50) | 2.32 | 265.20 | 0.824 | 0.574 |
| Class-Conditional ImageNet 512512 | ||||
| Model | FID | IS | Precision | Recall |
| BigGAN-deep [2] | 8.43 | 177.90 | 0.88 | 0.29 |
| StyleGAN-XL [67] | 2.41 | 267.75 | 0.77 | 0.52 |
| ADM [15] | 23.24 | 58.06 | 0.73 | 0.60 |
| ADM-U | 9.96 | 121.78 | 0.75 | 0.64 |
| ADM-G | 7.72 | 172.71 | 0.87 | 0.42 |
| ADM-G, ADM-U | 3.85 | 221.72 | 0.84 | 0.53 |
| Simple Diffusion (U-Net) [33] | 4.28 | 171.0 | - | - |
| Mask-GIT [5] | 7.32 | 156.0 | - | - |
| Simple Diffusion (U-ViT, L) | 4.53 | 205.3 | - | - |
| DiT-XL/2 [54] | 12.03 | 105.25 | 0.75 | 0.64 |
| DiT-XL/2-G (cfg=1.25) | 4.64 | 174.77 | 0.81 | 0.57 |
| DiT-XL/2-G (cfg=1.50) | 3.04 | 240.82 | 0.84 | 0.54 |
| SiT-XL [50] (cfg=1.50) | 2.62 | 252.21 | 0.84 | 0.57 |
| LiT-XL/2 | 14.00 | 92.84 | 0.76 | 0.62 |
| LiT-XL/2-G (cfg=1.50) | 3.69 | 207.97 | 0.85 | 0.53 |
As reported in IDDPM [53], a suitable variance can contribute to the initial steps of the diffusion process, thereby improving the log-likelihood. Accordingly, the authors set the variance learnable. Based on the finding, we propose to use the variance from the teacher diffusion model as a supervision for the student, but in a moderate way. The MSE between the reverse process variances of teacher and student can be calculated as:
where . With Eq. 5 and Eq. 6, we introduce a hybrid diffusion knowledge distillation objective, involving both the predicted noise and the variance:
where , are hyper-parameters. We thoroughly study some representative settings of our hybrid distillation objective as presented in Tab. 4, and identifies two key findings: (1) Compared to the baseline (distilling noise only), the proposed hybrid distillation contributes positively (50.79 vs. 51.13, measured by FID), but the variance distillation should be moderate. (e.g., set to 0.05). (2) Distilling only the variance is counterproductive for optimization, probably as the denoising capability is the major focus for diffusion.
Closing remarks.
So far, we have finalized the roadmap with 5 practical guidelines. We firmly answer how to convert a pre-trained DiT safely and efficiently to LiT while maintaining performance. Next, we will validate it on larger capacities (e.g., B/L/XL) and text-to-image generation.
4 Experiments
As this work focuses on safely and efficiently converting a pretrained baseline into a linear version without performance gap, we thus choose DiT [54] (for class-conditional image generation) and PixArt- [7] (for text-to-image generation) as baselines. The goal for LiT is to achieve comparable performance while maintaining an efficient training.
4.1 Class-Conditional Image Generation
256256 ImageNet.
We evaluate LiT on ImageNet [14] benchmark. LiT-S/2, B/2, L/2, XL/2 configurations are consistent with DiT, except that head numbers in linear attention are 2/3/4/4, following Guideline 2 in Sec. 3.2. DWC kernel size is set to 5 for all model variants. Detailed configuration are provided in Sec. A of the Appendix. We train 400K steps with a batch size of 256. For LiT-XL/2, we extend training to 1.4M steps (20 of DiT’s 7M steps). We used the AdamW [48] optimizer with a learning rate of and weight decay set to 0. We use 250 sampling steps, aligning with DiT. We initialize LiT’s parameters using the pre-trained DiT. , in hybrid knowledge distillation, are set to 0.5 and 0.05. Evaluation metrics include FID-50K [28], Inception Score [66] and Precision/Recall [40].
As reported in Tab. 5, LiT trained for only 100K steps already outperforms the DiT trained for 400K steps across different evaluation metrics and model variants. With extra training up to 400K steps, the model’s performance continues to improve. In Tab. 6, we compare LiT-XL/2 trained for 1.4M steps with high-performance baselines. Despite having only of the training steps of DiT-XL/2, LiT still competes on par with DiT (2.32 vs. 2.27). Besides, LiT competes favorably with several U-Net [64] based baselines. The results demonstrate that our LiT, when combined with suitable optimization strategies, can be reliably used in image generation applications.
512512 ImageNet.
We further validate LiT-XL/2 on the 512px setting. Using the pre-trained DiT-XL/2 [54] as the teacher, we initialize LiT-XL/2 with its weights. For knowledge distillation, we set and to 1.0 and 0.05, respectively, and train LiT-XL/2 for only 700K steps ( of DiT’s [54] 3M steps). Notably, unlike DiT, which used a batch size of 256, we adopt a smaller batch size of 128, requiring twice the training steps to complete 1 epoch (i.e., processing the entire training dataset once). Aside from this, all other hyper-parameter settings were consistent with the 256256 experiments. The results are reported in Tab. 7. Despite being trained for only 700K steps, which is equivalent in training epochs to 350K steps with a batch size of 256 on the same dataset, LiT, using pure linear attention, achieves an impressive FID of 3.69, comparable to DiT trained for 3M steps, reducing the training steps by . Additionally, LiT outperforms several strong baselines. These results demonstrate the effectiveness of our proposed 5 practical guidelines on higher-resolution setup.
4.2 Text-to-Image Generation
| Model | #Params. | Single | Two | Count. | Colors | Pos. | Attri. | Overall |
| LDM [63] | 1.4B | 0.92 | 0.29 | 0.23 | 0.70 | 0.02 | 0.05 | 0.37 |
| SDv1.5 [63] | 0.9B | 0.97 | 0.38 | 0.35 | 0.76 | 0.04 | 0.06 | 0.43 |
| SDv2.1 [63] | 0.9B | 0.98 | 0.51 | 0.44 | 0.85 | 0.07 | 0.17 | 0.50 |
| LlamaGen [69] | 0.8B | 0.71 | 0.34 | 0.21 | 0.58 | 0.07 | 0.04 | 0.32 |
| PixArt- [8] | 0.6B | 0.98 | 0.50 | 0.44 | 0.80 | 0.08 | 0.07 | 0.48 |
| PixArt- [7] | 0.6B | - | - | - | - | - | - | 0.52 |
| Lumina-Next [89] | 2.0B | - | - | - | - | - | - | 0.46 |
| SEED-X [18] | 17B | 0.97 | 0.58 | 0.26 | 0.80 | 0.19 | 0.14 | 0.49 |
| Chameleon [70] | 34B | - | - | - | - | - | - | 0.39 |
| LWM [45] | 7B | 0.93 | 0.41 | 0.46 | 0.79 | 0.09 | 0.15 | 0.47 |
| LiT (512px) | 0.6B | 0.97 | 0.43 | 0.42 | 0.79 | 0.09 | 0.12 | 0.47 |
| LiT (1024px) | 0.6B | 0.98 | 0.50 | 0.40 | 0.77 | 0.11 | 0.12 | 0.48 |
Experimental setup.
Building up with PixArt- [7] as the baseline, we develop LiT-0.6B with 2 linear attention heads and a DWC kernel size of 5. We leverage the pretrained SDXL [57] VAE Encoder and T5 [11] Encoder (i.e., Flan-T5-XXL) for image and text feature extraction, respectively. In 512px setting, we use PixArt- as the teacher to supervise the training of LiT-0.6B, with , set to 1.0, 0.05. As in Sec. 3.3, we inherit the weights of LiT from PixArt-, except for the parameters in the whole attention module. Subsequently, we train LiT-0.6B on an internal dataset, with a mixture of long (sharegpt caption, ) and short (original caption, ) captions. Token length is set to . We train with a learning rate of and batch size of 848 for 106K steps, which is efficient compared to the multi-stage training used for PixArt- [8]. This may also be related to our proposed weight inheritance strategy. We use 20 sampling steps, following PixArt-. In 1K setting, we basically follow the 512px seting for training and qualitative evaluation. We choose 512px LiT-0.6B as the starting point for training, with positional embeddings interpolation applied. We use the PixArt- [7] for 1K resolution as the teacher model. Starting from the 512px LiT-0.6B, we continue training on the internal datasets for 155K steps with a batch size of 248. We select suitable LiT-0.6B variants for sampling qualitative results and laptop deployment.
Results.
Tab. 8 compare the GenEval [19] results of LiT with those of other powerful methods. Despite relying solely on linear attention, LiT competes on par with its baseline PixArt-, and surpasses several strong text-to-image generation baselines, such as Lumina-Next [89], LlamaGen [69], and SDv1.5 [63]. Fig. 5, 6, and 7 (Sec. F of the Appendix) depict the 512px images generated by LiT. LiT-0.6B can produce exceptional, photorealistic images. Fig. 2 shows the sampled 1K resolution images generated by LiT. Our LiT-0.6B is able to generate images at a higher resolution with fine details and colorful contents, with pure attention in each block. These results show the generality of our proposed practical guidelines in text-to-image frameworks.
| Model | #Params. | Attention | Laptop | Latency (1K) | Latency (2K) |
| PixArt- [7] | 0.6B | KV Compress | ✗ | 4.38s | 32.16s |
| LiT | 0.6B | Linear | ✓ | 3.93s | 14.59s |
Efficiency Analysis
We compare the latency of LiT and our baseline PixArt- [7] for generating 1K and 2K resolution images, as shown in Tab. 9. The results are averaged over 10 runs on a 40GB NVIDIA A100 GPU. While PixArt- employs key-value compression [76] for attention acceleration, LiT’s linear attention proves more efficient, reducing latency by at 1K and at 2K resolution.
On-Device Offline Deployment
We deploy our LiT-0.6B for 1K resolution to a laptop powered by the Windows 11 operating system to validate its on-device capability. Specifically, we quantize the text encoder to 8 bits while using fp16 precision for other weights. The results (two photos) are provided in Sec. F of the Appendix. LiT-0.6B generates photorealistic 1K-resolution images offline, showcasing its efficiency for edge computing.
5 Conclusion and Limitation
This paper answers the question of how to safely and efficiently convert a pre-trained DiT for (text-to-)image generation to a linear DiT without compromising model capacity. We reexamine attention head number, weight inheritance, and knowledge distillation, and propose 5 useful guidelines as the response. In future work, we will focus on scalability and high-resolution (e.g., 4K) LiT generation.
Acknowledgement
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000.
Supplementary Material
Appendix A Model Configuration
We report the configurations of LiT variants in Tab. 10, which basically follow the hyper-parameters of DiT [54], except for using few heads. For class-conditional image generation, we set 2/3/4/4 heads for linear attention used in LiT. For LiT-XL/2, used for text-to-image tasks, we use 2 heads for linear attention.
| Model | Layers | Hidden size | Heads | Patch size |
| LiT-S | 12 | 384 | 2 | 2 |
| LiT-B | 12 | 768 | 3 | 2 |
| LiT-L | 24 | 1024 | 4 | 2 |
| LiT-XL | 28 | 1152 | 4 | 2 |
| LiT-XL⋄ | 28 | 1152 | 2 | 2 |
Appendix B Latency Analysis in Diffusion Transformer
We conduct a component-wise latency analysis of the Diffusion Transformer, with results shown in Fig. 7. The analysis was performed using the DiT-B/4 [54] model on an NVIDIA A100 GPU. The results indicate that the self-attention module accounts for 42.6% of the total latency of a DiT block. We attribute the observed considerable latency proportion to the quadratic computational complexity of the self-attention.
Appendix C Detailed Latency and Theoretical GMACs
We use one NVIDIA V100 GPU to evaluate the latency and theoretical GMACs of the DiT-S/B models with different numbers of heads. The task was to generate 256256 resolution images with a batch size of 8, following the experimental setup of Fig. 5 in the main paper (except for the GPU type), and the results are shown in Fig. 9.
We observe that both the small and base models on the V100 GPU exhibit a phenomenon similar to that on the A100 GPU: as theoretical GMACs increased, practical latency does not follow the same trend and even descend (in the B/2 model). This finding supports the generalizability of the free lunch effect of linear attention.
Appendix D Detailed Results on Attention Head Similarities
We take LiT-S/2 with 6 heads for visualization. Average cosine similarity among the attention maps of different heads is illustrated in Fig. 10.
Appendix E Detailed Results on Class-Conditional Image Generation
Detailed results on Sec. 4.
We provide detailed results for Tab. 1, Tab. 2, Tab. 3 and Tab. 4 in the main text in Tab. 11, Tab. 12, Tab. 13, and Tab. 14, respectively. In each table, we report results involving FID-50K [28] (without classifier-free guidance), Inception Score (IS) [66] and Precision/Recall [40]. IS, and Precision/Recall show results similar to FID-50K. As a result, the conclusions drawn in Sec. 4 of the main paper apply not only to metrics evaluating the distance between generated images and real images (e.g., FID-50K) but also to metrics reflecting the quality of the generated images themselves (e.g., IS).
Appendix F More Results on Text-to-Image Generation
We provide more text-to-image results in Fig. 13 and Fig. 14. As shown, LiT can accurately generate 512px photo-realistic images in various styles, themes, and content, whether the human instructions are simple or complicated. These results demonstrate that LiT effectively learns useful knowledge from the teacher model while maintaining exceptional computational efficiency, validating the effectiveness of our proposed cost-effective training strategy. Results of the offline laptop development are shown in Fig. 11.
Appendix G Pseudo-code
As presented in Alg. 1, we provide an example of the pseudo-code for the linear attention used in our LiT model for text-to-image generation for the laptop development. We set the kernel size of the depthwise convolution to 5.
Appendix H Full Related Work
Linear attention.
As a computationally efficient alternative to self-attention, linear attention [39] reduces computational complexity from quadratic to linear and has been proven effective in both visual understanding domain [3, 23, 24, 25, 1] and language domain [60, 80, 62, 61, 6]. EfficientViT [3] introduces a multi-scale linear attention with hardware-efficient operations to obtain a general vision backbone. Flatten Transformer [24] introduces focused linear attention to address the deficiencies in focus ability and feature diversity of linear attention, incorporating a focused function and depthwise convolution (DWC). SLAB [23] simplifies focused linear attention by retaining only the DWC component and introduces a progressive re-parameterized batch normalization to adapt offline batch normalization [36] for achieving low inference latency. These studies have been validated on visual perception tasks. Meanwhile, our work refines a linear attention module tailored for image generation tasks and identifies the free lunch of using few heads.
Efficient diffusion Transformer for image generation.
Limited by the quadratic computational complexity of self-attention, recent studies focus on developing efficient Transformer-style architectures for diffusion models. For example, DiM [71], ZigMa [35], and DiMSUM [56] explore implementing Mamba-based [22, 12] DiT-style [54] models. Diffusion-RMKV [17] studies RWKV-style [55] models for diffusion. Mediator [59] introduces an attention mediator to obtain an efficient diffusion Transformer with linear complexity. DiG [88] replaces the self-attention in DiT with gated linear attention to speed up training. LinFusion [47] and Sana [78] apply linear attention to U-Net-based [64] and Transformer-based [73] diffusion models, respectively, and train these models from scratch to generate high-quality images based on user instructions. Other studies [43, 42] explore efficient diffusion models through perspectives of low-bit quantization [26], feature map reusing [51, 41], and lightweight architecture design [87]. Differently, our work not only refines the design of linear attention but also introduces cost-effective training strategies, including weight inheritance and a novel hybrid diffusion distillation approach.
Advanced training method for diffusion models.
Some studies explores improved training strategies to enhance the optimization of diffusion models or achieve more efficient models. For example, CAN [4] introduces a condition-aware weight generation module to the diffusion Transformer, and demonstrate the technique can be further equipped with EffcientViT [3] to achieve both effectiveness and efficiency. REPA [84] proposes aligning the intermediate features of the denoising model with those extracted by a pre-trained visual encoder during the training of diffusion models. Some studies [65, 83, 82] explore distillation techniques to reduce the sampling steps of the diffusion model. Unlike the goal of reducing sampling steps, our proposed hybrid knowledge distillation focuses on an architectural perspective, exploring how a lightweight student model can learn from a computationally intensive teacher model.
| DiT | Attention | Resolution | Batch Size | Training Steps | FID-50K () | IS () | Precision () | Recall () |
| S/2 | softmax | 256 | 256 | 400K | 68.40 | - | ||
| S/2 | ReLU linear | 256 | 256 | 400K | 88.46 | 15.11 | 0.29 | 0.45 |
| S/2 | Simplified linear (ReLU) | 256 | 256 | 400K | 63.66 | 22.16 | 0.38 | 0.58 |
| S/2 | focused linear (ReLU) | 256 | 256 | 400K | 63.05 | 22.49 | 0.39 | 0.58 |
| S/2 | focused linear (GELU) | 256 | 256 | 400K | 70.83 | 19.41 | 0.36 | 0.54 |
| B/2 | softmax | 256 | 256 | 400K | 43.47 | - | ||
| B/2 | ReLU linear | 256 | 256 | 400K | 56.92 | 25.80 | 0.42 | 0.59 |
| B/2 | Simplified linear (ReLU) | 256 | 256 | 400K | 42.11 | 34.60 | 0.48 | 0.63 |
| B/2 | focused linear (ReLU) | 256 | 256 | 400K | 40.58 | 35.98 | 0.50 | 0.63 |
| B/2 | focused linear (GELU) | 256 | 256 | 400K | 58.86 | 24.23 | 0.42 | 0.57 |
| DiT | Head | Resolution | Batch Size | Training Steps | FID-50K () | IS () | Precision () | Recall () |
| S/2 | 1 | 256 | 256 | 400K | 64.42 | 21.54 | 0.380 | 0.574 |
| S/2 | 2 | 256 | 256 | 400K | 63.24 | 22.07 | 0.385 | 0.570 |
| S/2 | 3 | 256 | 256 | 400K | 63.21 | 22.08 | 0.386 | 0.583 |
| S/2 | 6 | 256 | 256 | 400K | 63.66 | 22.16 | 0.383 | 0.580 |
| S/2 | 48 | 256 | 256 | 400K | 78.76 | 17.46 | 0.322 | 0.482 |
| S/2 | 96 | 256 | 256 | 400K | 116.00 | 11.49 | 0.224 | 0.261 |
| B/2 | 1 | 256 | 256 | 400K | 41.77 | 34.78 | 0.487 | 0.631 |
| B/2 | 2 | 256 | 256 | 400K | 41.39 | 35.59 | 0.494 | 0.631 |
| B/2 | 3 | 256 | 256 | 400K | 40.86 | 35.79 | 0.497 | 0.629 |
| B/2 | 12 | 256 | 256 | 400K | 42.11 | 34.60 | 0.484 | 0.631 |
| B/2 | 96 | 256 | 256 | 400K | 68.30 | 20.45 | 0.375 | 0.531 |
| B/2 | 192 | 256 | 256 | 400K | 112.39 | 12.07 | 0.240 | 0.282 |
| L/2 | 1 | 256 | 256 | 400K | 24.46 | 57.36 | 0.600 | 0.637 |
| L/2 | 2 | 256 | 256 | 400K | 24.37 | 57.02 | 0.599 | 0.622 |
| L/2 | 4 | 256 | 256 | 400K | 24.04 | 59.02 | 0.597 | 0.636 |
| L/2 | 16 | 256 | 256 | 400K | 25.25 | 54.67 | 0.587 | 0.632 |
| XL/2 | 1 | 256 | 256 | 400K | 21.13 | 65.06 | 0.619 | 0.632 |
| XL/2 | 2 | 256 | 256 | 400K | 20.66 | 65.39 | 0.624 | 0.636 |
| XL/2 | 4 | 256 | 256 | 400K | 20.82 | 65.52 | 0.619 | 0.632 |
| XL/2 | 16 | 256 | 256 | 400K | 21.69 | 63.06 | 0.617 | 0.628 |
| Load | Iterations | FFN | Modulation Layer | Attention | FID-50K () | IS () | Precision () | Recall () |
| model | 400K | ✓ | ✓ | ✗ | 56.07 | 25.62 | 0.418 | 0.608 |
| ema | 400K | ✓ | ✓ | ✗ | 56.07 | 25.61 | 0.416 | 0.601 |
| model | 200K | ✓ | ✓ | ✗ | 57.84 | 24.72 | 0.408 | 0.600 |
| model | 300K | ✓ | ✓ | ✗ | 56.95 | 25.04 | 0.414 | 0.608 |
| model | 400K | ✓ | ✓ | ✗ | 56.07 | 25.62 | 0.418 | 0.608 |
| model | 600K | ✓ | ✓ | ✗ | 54.80 | 26.65 | 0.424 | 0.613 |
| model | 800K | ✓ | ✓ | ✗ | 53.83 | 27.16 | 0.425 | 0.614 |
| model | 600K | ✓ | ✓ | Q, K, V | 55.29 | 26.09 | 0.419 | 0.619 |
| model | 600K | ✓ | ✓ | K, V | 55.07 | 26.38 | 0.422 | 0.609 |
| model | 600K | ✓ | ✓ | V | 54.93 | 26.44 | 0.427 | 0.612 |
| model | 600K | ✓ | ✓ | Q | 54.82 | 26.72 | 0.423 | 0.605 |
| model | 600K | ✓ | ✓ | O | 54.84 | 26.33 | 0.425 | 0.607 |
| Model Size | Iterations | Teacher | Training Steps | FID-50K () | IS () | Precision () | Recall () | ||
| S/2 | 800K | DiT-S/2 | 0.1 | 0.0 | 400K | 55.11 | 26.28 | 0.419 | 0.614 |
| S/2 | 800K | DiT-XL/2 | 0.0 | 0.0 | 400K | 53.83 | 27.16 | 0.425 | 0.614 |
| S/2 | 800K | DiT-XL/2 | 0.1 | 0.0 | 400K | 53.05 | 27.43 | 0.431 | 0.609 |
| S/2 | 800K | DiT-XL/2 | 0.05 | 0.0 | 400K | 53.41 | 27.26 | 0.427 | 0.610 |
| S/2 | 800K | DiT-XL/2 | 0.5 | 0.0 | 400K | 51.13 | 28.89 | 0.438 | 0.616 |
| S/2 | 800K | DiT-XL/2 | 0.1 | 0.05 | 400K | 52.76 | 27.70 | 0.431 | 0.620 |
| S/2 | 800K | DiT-XL/2 | 0.0 | 0.05 | 400K | 53.49 | 27.26 | 0.429 | 0.609 |
| S/2 | 800K | DiT-XL/2 | 0.05 | 0.05 | 400K | 53.14 | 27.46 | 0.431 | 0.609 |
| S/2 | 800K | DiT-XL/2 | 0.5 | 0.05 | 400K | 50.79 | 29.17 | 0.443 | 0.618 |